Analyse the cross price elasticity of demand for wheat and rice when a change in the

price of wheat from Rs.70 to Rs. 90 results in a change in the quantity demand for

wheat from kg 3000 to kg 5000 in the market. Interpret the value of the coefficient.


1
Expert's answer
2021-04-23T12:08:27-0400

"E_d= \frac{\frac{\Delta Q}{Q}}{\frac{\Delta P}{P}} = \frac{\frac{2000}{3000}}{\frac{20}{70}} = \frac{7}{3} = 2,33"

A postive cross price elasticity means that both products are substistutive. That means increase in price of one of them will increase the quantity demanded of the other one